# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v6

@@ 1 Introduction
However, most of the existing work on GS (see (?) for an overview) has focused on formalisms based on abstract arguments (?), treated as atomic entities without internal structure. This abstraction, while useful in many contexts, fails to capture the rich logical structure that is often present in real-world arguments. Meanwhile, various forms of structured argumentation frameworks (see (?) for an overview) allow to represent the internal composition of arguments and the relationships between their components, and thus offer a more expressive and realistic model in agent-based domains, such as user modeling  (?), scientific debates (?), and model reconciliation settings (?). However, this greater reliance on the logical (e.g., background or contextual) knowledge can lead to more problematic cases where this information is incomplete. Yet, it is only recently that the study of GS for structured argumentation has received attention, e.g., for deductive argumentation (?) by Heyninck et al. (?), for assumption-based argumentation (ABA) (?) by Skiba et al. (?), for ASPIC+ (?) by Spaans (?) and Prakken (?), and for a restricted type of structured argumentation in the form of statement graphs (SGs) (?) by Jedwabny et al. (?).
@@
We selected SGs as our targeted form of structured argumentation in this paper. We chose this starting point for our analysis, rather than other forms of structured argumentation, because they naturally accommodate attack and support relations and their GS are thus naturally relatable to those of QBAFs, which are the most widely studied form of argumentation when it comes to GS (?). This allows us to undertake an interesting analysis of properties of our GS in direct comparison with those for QBAFs, given their widespread application.
@@
For an illustration of the need for GS in structured argumentation, Figure 1 shows a debate about climate change as an SG. SGs use a restricted type of structured arguments in the form of statements, each consisting of a premise (a compound of literals, which may possibly be $\top$, i.e., true) and a claim (a literal). A typical statement ($\alpha_{1}$) may combine evidence about rising CO2 emissions ($a$) and increasing global temperatures ($b$), given that they are both supported by facts ($\alpha_{2}$ and $\alpha_{3}$), to support the claim that climate change is human-caused ($c$). Such an SG may be the result of applying argument mining to text, e.g., as overviewed in  (?). Such argument mining techniques coupled with suitable GS for SGs could allow for better more expressive representation of arguments while handling incomplete information in the modelling of debates, a rapidly growing research area which utilises many forms of argumentation (??????), especially given the groundbreaking potential of large language models (?).
@@
Figure 1: Example SG (including arbitrary weights $\tau$, see §3). Green edges with a plus denote support; red edges with a minus denote attack. The SG includes incomplete information about the statement $\alpha_{4}$, since its premise $d$ is not grounded in facts. However, the given weight for $\alpha_{4}$ provides information that our GS can use.
@@
The GS proposed by Jedwabny et al. (?) for SGs uses T-norms and T-conorms from fuzzy logic to aggregate the strengths of a statement’s components, but this GS faces challenges when dealing with incomplete information since it relies on an assumption of complete knowledge and presence of a statement’s supporting elements. This requirement can limit its applicability in real-world scenarios, where such information is almost always incomplete. For example, in Figure 1, the strength of the statement $\alpha_{1}$ is reduced (e.g., in the eyes of a modelled audience) by a statement ($\alpha_{4}$) questioning the reliability of CO2 measurements, without being evidenced (there is no support for $d$). To date, existing approaches to GS in structured argumentation struggle to handle such incompleteness, potentially leading to oversimplified or incorrect conclusions, and thus poor modelling of debates.
@@
To address this limitation, we introduce a novel methodology for GS in SGs that can naturally accommodate complete and incomplete information. Our approach is modular in that it allows for different GS to be instantiated based on the requirements of a given task and context. Specifically, our GS can be instantiated with any existing GS for QBAFs. This existing GS is used to dialectically evaluate the literals in a premise, based on their attacking and supporting evidence, before these “intermediate” evaluations are aggregated (depending on the statement’s construction) to give a final strength for the statement.
@@
  • Using these properties, we conduct a comprehensive theoretical analysis of our two proposed instantiations, two instantiations of the T-norm semantics (?) and two existing GS for QBAFs, demonstrating the strengths and limitations of each approach.
@@ 2 Related Work
GS have received significant attention within abstract argumentation in recent years, given their professed alignment with human reasoning (??) and applicability to real-world settings (??). This has given rise to a rich field of research, with a broad repertoire of GS (???), in-depth analyses of GS’ behaviour (????), and a number of open-source computational toolsets (??). One issue in this area is that many abstract GS, given their recursive nature, do not satisfy the uniqueness property in assigning strengths in cyclic graphs (??). This must be considered if our acyclicity restriction is relaxed in future work.
With regards to other approaches to GS in structured argumentation, our approach differs from existing work as follows. Within ASPIC+, Spaans (?) introduced initial strength functions, with corresponding properties, though they use weights on rules and do not consider explicit support relations. Further, Prakken (?) gives a formal model of argument strength and deploys it. However, the main difference between this work and our approach is our leveraging of existing bipolar abstract GS for handling incomplete information. We leave to future work the assessment of this approach with respect to our proposed properties. Also for ASPIC+, extension-based semantics have been used for cases with incomplete information (?).
@@
Finally, a related but distinct subfield is probabilistic argumentation (???????), but in all of these works, either support relations are not considered or strengths are associated with probabilities and/or are not assigned to individual arguments.
@@ 3 Preliminaries
We consider a propositional language $\mathcal{L}$ comprising a finite set of atoms, including a special atom $\top$ to represent truth. A literal is an atom $\phi$ or its negation $\neg\phi$. Note that ¬($\neg(\neg\phi)\equiv\phi.$ A compound is an expression constructed from literals using connectives $\wedge$ and $\vee$.
@@
A statement attacks or supports another if the former gives justification against or for (respectively) the latter’s premise:
@@
A GS $\sigma:\mathcal{X}\rightarrow[0,1]$ assigns each statement $\alpha\in\mathcal{X}$ a strength $\sigma^{\mathcal{G}}(\alpha)\in[0,1]$. [^5] When defining GS generally, we drop the $\mathcal{G}$ superscript.
In essence, GS offer a principled approach to evaluating statements quantitatively. We refer informally to any GS that takes into account the premises of any statement as a structured GS, and any that does not as an abstract GS. It should be noted that while abstract GS can be applied to SGs, structured GS cannot be applied to frameworks using abstract argumentation such as QBAFs, since the statements therein are abstract entities.
@@
Next, we recall two notions from (?). [^6] We omit the acyclicity condition since we use acyclic graphs.
@@
A triple (,,¬)tensor-productdirect-sum($\otimes$,$\oplus$,$\neg$)( ⊗ , ⊕ , ¬ ), where $\otimes$ is a T-norm, $\oplus$ is a T-conorm, and $\neg$ is a negation, is called a De Morgan triple iff $\forall v_{1},v_{2}\in[0,1]$, $\otimes(v_{1},v_{2})=\neg(\oplus(\neg(v_{1}),\neg(v_{2}))$, and $\oplus(v_{1},v_{2})=\neg(\otimes(\neg(v_{1}),\neg(v_{2}))$.
@@
Given a De Morgan triple (,,¬)tensor-productdirect-sum($\otimes$,$\oplus$,$\neg$)( ⊗ , ⊕ , ¬ ), the T-norm semantics $\sigma_{T}$ is a structured GS s.t. for any $\alpha\in\mathcal{X}$, where $\sigma^{\mathcal{G}}_{T}(\alpha)=\bigoplus_{T\in\mathcal{T}^{\mathcal{G}}(\alpha)}\mathcal{O}(T)$, $\mathcal{O}(T)=\mathcal{I}(T)\otimes\neg\bigoplus_{T^{\prime}\in\mathcal{T}^{\mathcal{G}}(\alpha^{\prime}\in\mathcal{X})\text{, }T^{\prime}\text{ attacks }T}\mathcal{I}(T^{\prime})$ and $\mathcal{I}(T)=\bigotimes_{\alpha\in T}\tau(\alpha)$.
While the T-norm semantics is modular wrt the De Morgan triple, the methodology is based on CSTs, and thus it can be seen that it inherently relies on complete information.
@@
  • $\Sigma$ is s.t. $\Sigma(())=0$, where ()()( ) is an empty sequence, and, for $v_{1},\ldots,v_{n}\in[0,1]$ ($n\geq 1$), if $n=1$, then $\Sigma((v_{1}))=v_{1}$; if $n=2$, then $\Sigma((v_{1},v_{2}))=v_{1}+v_{2}-v_{1}\cdot v_{2}$; and if $n>2$, then $\Sigma((v_{1},\ldots,v_{n}))=\Sigma(\Sigma((v_{1},\ldots,v_{n-1})),v_{n})$;
@@ 4 Gradual Semantics for Statement Graphs
Note that partially-complete and incomplete statements are in the same spirit as potential arguments in ABA (?), but the latter are used for computational purposes, as steps towards complete statements, rather than as representations in their own right, as we do.
@@
Consider an SG, as illustrated in Figure 2, in which $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$, $\alpha_{1}=\langle b,a\rangle$, $\alpha_{2}=\langle c,b\rangle$ and $\alpha_{3}=\langle d,b\rangle$, giving $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1})\}$ and $\mathcal{A}=\emptyset$ (the $\tau$ values are irrelevant here). In this first case, $\alpha_{1}$ incomplete as it has no CSTs. If we add the statement $\alpha_{4}=\langle\top,c\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2})\}$, $\alpha_{1}$ becomes partially-complete as it has one CST, but $\alpha_{3}$, which has a path to $\alpha_{1}$, does not have a CST. Finally, if we add the statement $\alpha_{5}=\langle\top,d\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2}),(\alpha_{5},\alpha_{3})\}$, $\alpha_{1}$ becomes complete as all statements with a path to it have CSTs. Crucial here is the fact that in the incomplete case the T-norm semantics ignores the (potentially useful) information from both $\alpha_{2}$ and $\alpha_{3}$, regardless of $\tau$, while in the partially-complete case it ignores that of $\alpha_{3}$.
Figure 2: SG from Example 1 with initial (added) statements shown in solid (striped, respectively) turquoise, no attacks, initial (added) supports as solid (dashed, respectively) green arrows and some arbitrary $\tau$ (not shown here). The statement $\alpha_{1}\in\mathcal{X}$ is incomplete when $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$, but is partially-complete when $\alpha_{4}$ (or $\alpha_{5}$) is added and complete when both $\alpha_{4}$ and $\alpha_{5}$ are added.
@@ 4.1 Methodology
Note that conjunctive premises represent a challenge for handling incomplete information, as the failure of any single premise invalidates the entire statement. For that matter, the DC semantics has been designed such that certain desirable theoretical properties will be satisfied, as we will see in §4.2. Importantly, the DC semantics represents a natural starting point for our methodology because conjunctive premises are typically abound in real-world argumentative reasoning. For instance, when people construct arguments, they typically combine multiple pieces of evidence conjunctively to support their claims. This pattern is evident in scientific argumentation, legal reasoning and everyday discourse, where people naturally aggregate supporting facts with “and” connectives (?). In future work, modular structured GS could be instantiated for different structures of SGs, e.g., for SGs with disjunctions as premises, a probabilistic sum could be used for the premise aggregation function. This would align with the intuition that multiple independent supporting lines of reasoning may strengthen a statement’s overall strength. For other logical structures such as statements in conjunctive normal form, a hierarchical aggregation approach could be employed as explored, i.e., first combining literals within clauses using disjunctive operators, then aggregating clauses conjunctively.
@@ 4.2 Properties
The first property (?) is a fundamental requirement for GS, stating that a statement’s strength only depends on the statements that are connected to it.
@@
Given a second SG $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, where $\alpha_{2}=\alpha_{1}$ or $\alpha_{3}=\alpha_{1}$ $\forall(\alpha_{2},\alpha_{3})\in(\mathcal{A}^{\prime}\cup\mathcal{S}^{\prime})\setminus(\mathcal{A}\cup\mathcal{S})$, a GS $\sigma$ satisfies directionality iff for any $\alpha_{2}\in\mathcal{X}$, where there exists no path from $\alpha_{1}$ to $\alpha_{2}$ via $\mathcal{A}\cup\mathcal{S}$ and $\tau^{\prime}(\alpha_{3})=\tau(\alpha_{3})$ $\forall\alpha_{3}\in\mathcal{X}$, it holds that $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Next, Jedwabny et al. (?) introduced a property which requires that the length of a proof in the logic should not affect the strength of a statement.
@@
This property is indeed intuitive in a number of settings where we would prioritise the logical consistency over approximate evaluations of statements. However, in some settings, e.g., when modelling human discourse, it may be the case that the same logical conclusions, when reached over the course of multiple statements, are weaker, or even stronger, than those stated succinctly (e.g., see the study of priority in (?)). For example, given the statements instantiated in rewriting, consider the effect of adding an statement $\alpha_{4}=\langle\Phi_{4},\Psi_{4}\rangle$, where $\Psi\equiv\neg x_{1}$. In some settings may be desirable that the weakening effect of $\alpha_{4}$ on $\alpha_{3}$ is lesser than that on $\alpha_{1}$ due to the distance. We leave a formalisation of this (non-trivial) property to future work. We would, however, argue that rewriting may not be universally applicable.
@@
Continuing from Example 2, let us introduce two additional statements: $\alpha_{5}=\langle a,e\rangle$ (“CO2 emissions are rising, therefore extreme weather events are increasing”), and $\alpha_{6}=\langle e\wedge b,c\rangle$ (“Extreme weather events and global temperatures are increasing, therefore climate change is human-caused”). Let $\tau(\alpha_{5})=1$ and $\tau(\alpha_{6})=0.8$. $\mathcal{A}=\{(\alpha_{4},\alpha_{1}),(\alpha_{4},\alpha_{5})\}$. Under the DC semantics, we get $\sigma_{\wedge_{D}}^{\mathcal{G}}(\alpha_{1})=0.877$ and $\sigma_{\wedge_{D}}^{\mathcal{G}}(\alpha_{6})=\sigma_{D}^{\mathcal{G}^{1}}(e)\times\sigma_{D}^{\mathcal{G}^{1}}(b)=1\times 0.957=0.957$. In contrast, under T-norm-p semantics, we get $\sigma_{T_{p}}^{\mathcal{G}}(\alpha_{1})\!=\!\sigma_{T_{p}}^{\mathcal{G}}(\alpha_{6})\!=\!0.432$. We can see that the DC semantics violate the rewriting property, as $\alpha_{1}$ and $\alpha_{6}$ have different strengths despite making the same claim through different reasoning paths. The T-norm-p semantics, however, is agnostic to the different lengths of the reasoning chain.
Next, Jedwabny et al. (?) state that a statement without at least one supporter for each of its premises is bottom-strength.
@@
This property is clearly useful in cases under complete information, i.e., in the first case in Definition 9, but in the other two cases, we posit that this property may be too demanding.
@@
In Example 3, weak provability would not require that $\alpha_{4}$’s strength is zero, allowing a form of “trust” in the statement’s weight in the absence of complete information providing the logical reasoning to support its premises. This means the information is not lost in the partially-complete or incomplete cases, unless it is assigned the minimal strength by the GS.
@@
The desirability of this property is clear, demonstrating an intuitive starting point for the strength of a statement, regardless of its completeness from Definition 9.
@@
First, we consider the effect of adding a statement that has been assigned a zero strength, i.e., that it should not affect the strength of any statement it attacks or supports, in line with neutrality (?) for abstract GS.

### v7

@@ 1 Introduction
However, most of the existing work on GS (see (?) for an overview) has focused on formalisms based on abstract arguments (?), treated as atomic entities without internal structure. This abstraction, while useful in many contexts, fails to capture the rich logical structure that is often present in real-world arguments. Meanwhile, various forms of structured argumentation frameworks (see (?) for an overview) allow to represent the internal composition of arguments and the relationships between their components. They thus offer a more expressive and realistic model in agent-based domains, such as user modeling  (?), scientific debates (?), and model reconciliation settings (?). However, this greater reliance on the logical (e.g., background or contextual) knowledge can lead to more problematic cases where this information is incomplete, where the completeness of a statement (as we formally define in §4) refers to its grounding in facts. Yet, it is only recently that the study of GS for structured argumentation has received attention, e.g., for deductive argumentation (?) by Heyninck et al. (?), for assumption-based argumentation (ABA) (?) by Skiba et al. (?), for ASPIC+ (?) by Spaans (?) and Prakken (?), and for a restricted type of structured argumentation in the form of statement graphs (SGs) (?) by Jedwabny et al. (?).
@@
We selected SGs as our targeted form of structured argumentation in this paper. We chose this starting point for our analysis, rather than other forms of structured argumentation, because they naturally accommodate attack and support relations. Their GS are thus naturally relatable to those of QBAFs, which are the most widely studied form of argumentation when it comes to GS (?). This allows us to undertake an interesting analysis of properties of our GS in direct comparison with those for QBAFs, given their widespread application.
@@
To illustrate the need for GS in structured argumentation, Figure 1 shows a debate about climate change represented as an SG. SGs use a restricted type of structured arguments in the form of statements, each consisting of a premise (a compound of literals, which may possibly be $\top$, i.e., true) and a claim (a literal). A typical statement ($\alpha_{1}$) may combine evidence about rising CO2 emissions ($a$) and increasing global temperatures ($b$), given that they are both supported by facts ($\alpha_{2}$ and $\alpha_{3}$), to support the claim that climate change is human-caused ($c$). Such an SG may be the result of mining arguments from text, e.g., as overviewed in  (?). Such argument mining techniques coupled with suitable GS for SGs could allow for more expressive representation of arguments while handling incomplete information in the modelling of debates, a rapidly growing research area utilising many forms of argumentation (??????), especially given the groundbreaking potential of large language models (?).
@@
Figure 1: Example SG (with arbitrary weights $\tau$, see §3). Green edges with a plus denote support; red edges with a minus denote attack. The SG includes incomplete information about the statement $\alpha_{4}$, since its premise $d$ is not grounded in facts. However, the given weight for $\alpha_{4}$ provides information that our GS can use.
@@
The GS proposed by Jedwabny et al. (?) for SGs uses T-norms and T-conorms from fuzzy logic to aggregate the strengths of a statement’s components, but this GS faces challenges when dealing with incomplete information as it relies on complete knowledge and the presence of a statement’s supporting elements. This requirement can limit its applicability in real-world scenarios, where such information is almost always incomplete. For example, in Figure 1, the strength of the statement $\alpha_{1}$ is reduced (e.g., in the eyes of a modelled audience) by a statement ($\alpha_{4}$) questioning the reliability of CO2 measurements, without being evidenced (there is no support for $d$). To date, existing approaches to GS in structured argumentation struggle to handle such incompleteness, potentially leading to oversimplified or incorrect conclusions, and thus poor modelling of debates.
@@
To address this limitation, we introduce a novel methodology for GS in SGs that can naturally accommodate complete and incomplete information. Our approach is modular in that it allows for different GS to be instantiated based on the requirements of a given task and context. Specifically, our methodology for GS can be instantiated with any existing GS for QBAFs. This existing GS is used to dialectically evaluate the literals in a premise, based on their attacking and supporting evidence, before these “intermediate” evaluations are aggregated (depending on the statement’s construction) to give a final strength for the statement.
@@
  • Using these properties, we conduct a comprehensive theoretical analysis of our two proposed instantiations, two instantiations of the T-norm semantics (?) and two existing GS for QBAFs, discussing the strengths and limitations of each approach.
@@ 2 Related Work
GS have received significant attention within abstract argumentation in recent years, given their professed alignment with human reasoning (??) and applicability to real-world settings (??). This has given rise to a rich field of research, with a broad repertoire of GS (???), in-depth analyses of GS’ behaviour (????), and a number of open-source computational toolsets (??). One issue in this area is that many abstract GS, e.g., those we use (??), do not satisfy the uniqueness property in assigning strengths in cyclic graphs due to their recursive nature (??). This must be considered if our acyclicity restriction is relaxed in future work.
There have been other works related to GS in structured argumentation, differing from our work as follows. Within ASPIC+, Spaans (?) introduced initial strength functions, with corresponding properties, though they use weights on rules and do not consider explicit support relations. Further, Prakken (?) gives a formal model of argument strength across graphs and deploys it. A major difference between this work and our approach is our leveraging of existing bipolar abstract GS for handling incomplete information. We leave to future work the assessment of this approach with respect to our proposed properties. Also for ASPIC+, extension-based semantics have been used for cases with incomplete information (?).
@@
Finally, probabilistic argumentation (??????????), is a related but distinct subfield where strengths, which may be over (sets of) arguments, relations or both, must adhere to strict probabilistic principles, whereas GS may not.
@@ 3 Preliminaries
We consider a propositional language $\mathcal{L}$ comprising a finite set of atoms, including a special atom $\top$ to represent truth. A literal is an atom $\phi$ or its negation $\neg\phi$. Note that $\neg(\neg\phi)\equiv\phi.$ A compound is an expression, limited in this paper to be constructed from literals using the connective $\wedge$.
@@
A statement attacks or supports another if the former gives justification against or for the latter’s premise:
@@
A GS $\sigma:\mathcal{X}\rightarrow[0,1]$ assigns each statement $\alpha\in\mathcal{X}$ a strength $\sigma^{\mathcal{G}}(\alpha)\in[0,1]$. [^5] With an abuse of notation, when referring to GS generally, we drop the $\mathcal{G}$ superscript for clarity.
In essence, GS offer a principled approach to evaluating statements quantitatively. We refer informally throughout the paper to any GS that takes into account the premises of any statement as a structured GS, and any that does not as an abstract GS. It should be noted that while abstract GS can be applied to SGs, structured GS cannot be applied to frameworks using abstract argumentation such as QBAFs, since the statements therein are abstract entities.
@@
Next, we recall two notions from (?). [^6] We omit the acyclicity condition since we use acyclic graphs. The first gives the notion indicating when an argument’s premises is grounded in facts, while the second defines operators used in the GS.
@@
A triple $(\otimes,\oplus,\neg)$, where $\otimes$ is a T-norm, $\oplus$ is a T-conorm, and $\neg$ is a negation, is called a De Morgan triple iff $\forall v_{1},v_{2}\in[0,1]$, $\otimes(v_{1},v_{2})=\neg(\oplus(\neg(v_{1}),\neg(v_{2}))$, and $\oplus(v_{1},v_{2})=\neg(\otimes(\neg(v_{1}),\neg(v_{2}))$.
@@
Given a De Morgan triple $(\otimes,\oplus,\neg)$, the T-norm semantics $\sigma_{T}$ is a structured GS s.t. for any $\alpha\in\mathcal{X}$, where $\sigma^{\mathcal{G}}_{T}(\alpha)=\bigoplus_{T\in\mathcal{T}^{\mathcal{G}}(\alpha)}\mathcal{O}(T)$, $\mathcal{O}(T)=\mathcal{I}(T)\otimes\neg\bigoplus_{T^{\prime}\in\mathcal{T}^{\mathcal{G}}(\alpha^{\prime}\in\mathcal{X})\text{, }T^{\prime}\text{ attacks }T}\mathcal{I}(T^{\prime})$ and $\mathcal{I}(T)=\bigotimes_{\alpha\in T}\tau(\alpha)$.
While the T-norm semantics is modular wrt the De Morgan triple, the methodology is based on a statement’s CSTs, and thus it can be seen that it inherently relies on complete information, i.e., a statement’s grounding in facts.
@@
  • $\Sigma$ is s.t. $\Sigma(())=0$, where $()$ is an empty sequence, and, for $v_{1},\ldots,v_{n}\in[0,1]$ ($n\geq 1$), if $n=1$, then $\Sigma((v_{1}))=v_{1}$; if $n=2$, then $\Sigma((v_{1},v_{2}))=v_{1}+v_{2}-v_{1}\cdot v_{2}$; and if $n>2$, then $\Sigma((v_{1},\ldots,v_{n}))=\Sigma(\Sigma((v_{1},\ldots,v_{n-1})),v_{n})$;
@@ 4 Gradual Semantics for Statement Graphs
Note that partially-complete and incomplete statements are in the same spirit as potential arguments in ABA (?), but the latter are used for computational purposes, as steps towards complete statements, rather than as representations in their own right, as we do. Further investigation, e.g. on the equivalent in ASPIC+, is left to future work.
@@
Consider the SG in Figure 2, in which $\mathcal{X}=\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}$, $\alpha^{\prime}_{1}=\langle b,a\rangle$, $\alpha^{\prime}_{2}=\langle c,b\rangle$ and $\alpha^{\prime}_{3}=\langle d,b\rangle$, giving $\mathcal{S}=\{(\alpha^{\prime}_{2},\alpha^{\prime}_{1}),(\alpha^{\prime}_{3},\alpha^{\prime}_{1})\}$ and $\mathcal{A}=\emptyset$ (the $\tau$ values are irrelevant here). In this first case, $\alpha^{\prime}_{1}$ incomplete as it has no CSTs. If we add the statement $\alpha^{\prime}_{4}=\langle\top,c\rangle$, leading to $\mathcal{S}=\{(\alpha^{\prime}_{2},\alpha^{\prime}_{1}),(\alpha^{\prime}_{3},\alpha^{\prime}_{1}),(\alpha^{\prime}_{4},\alpha^{\prime}_{2})\}$, $\alpha^{\prime}_{1}$ becomes partially-complete as it has one CST, but $\alpha^{\prime}_{3}$, which has a path to $\alpha^{\prime}_{1}$, does not have a CST. Finally, if we add the statement $\alpha^{\prime}_{5}=\langle\top,d\rangle$, leading to $\mathcal{S}=\{(\alpha^{\prime}_{2},\alpha^{\prime}_{1}),(\alpha^{\prime}_{3},\alpha^{\prime}_{1}),(\alpha^{\prime}_{4},\alpha^{\prime}_{2}),(\alpha^{\prime}_{5},\alpha^{\prime}_{3})\}$, $\alpha^{\prime}_{1}$ becomes complete as all statements with a path to it have CSTs. Crucial here is the fact that in the incomplete case the T-norm semantics ignores the (potentially useful) information from both $\alpha^{\prime}_{2}$ and $\alpha^{\prime}_{3}$, regardless of $\tau$, while in the partially-complete case it ignores that of $\alpha^{\prime}_{3}$.
Figure 2: SG from Example 1 with initial (added) statements shown in solid (striped, respectively) turquoise, no attacks, initial (added) supports as solid (dashed, respectively) green arrows and some arbitrary $\tau$ (not shown here). The statement $\alpha^{\prime}_{1}\in\mathcal{X}$ is incomplete when $\mathcal{X}=\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}$, but is partially-complete when $\alpha^{\prime}_{4}$ (or $\alpha^{\prime}_{5}$) is added and complete when both $\alpha^{\prime}_{4}$ and $\alpha^{\prime}_{5}$ are added.
@@ 4.1 Methodology
Note that conjunctive premises represent a challenge for handling incomplete information, as the failure of any single premise invalidates the entire statement. For that matter, the DC semantics has been designed such that certain desirable theoretical properties will be satisfied, as we will see in §4.2. Importantly, the DC semantics represents a natural starting point for our methodology because conjunctive premises are typically abound in real-world argumentative reasoning. For instance, when people construct arguments, they typically combine multiple pieces of evidence conjunctively to support their claims. This pattern is evident in scientific argumentation, legal reasoning and everyday discourse, where people naturally aggregate supporting facts with “and” connectives (?). In future work, modular structured GS could be instantiated for different structures of SGs, e.g., for SGs with disjunctions as premises, a probabilistic sum could be used for the premise aggregation function. This would align with the intuition that multiple independent supporting lines of reasoning may strengthen a statement’s overall strength. For other logical structures such as statements in conjunctive normal form, a hierarchical aggregation approach could be employed, i.e., first combining literals within clauses using disjunctive operators, then aggregating clauses conjunctively.
@@ 4.2 Properties
The first property (?) is a fundamental requirement for GS, stating that a statement’s strength only depends on statements that are connected to it, and are thus relevant to it.
@@
Given two SGs $\mathcal{G}$ and $\mathcal{G}^{\prime}=\langle\mathcal{X}\cup\{\alpha_{1}\},\mathcal{A}^{\prime},\mathcal{S}^{\prime},\tau^{\prime}\rangle$, where $\alpha_{2}=\alpha_{1}$ or $\alpha_{3}=\alpha_{1}$ $\forall(\alpha_{2},\alpha_{3})\in(\mathcal{A}^{\prime}\cup\mathcal{S}^{\prime})\setminus(\mathcal{A}\cup\mathcal{S})$, a GS $\sigma$ satisfies directionality iff for any $\alpha_{2}\in\mathcal{X}$, where there exists no path from $\alpha_{1}$ to $\alpha_{2}$ via $\mathcal{A}\cup\mathcal{S}$ and $\tau^{\prime}(\alpha_{3})=\tau(\alpha_{3})$ $\forall\alpha_{3}\in\mathcal{X}$, it holds that $\sigma^{\mathcal{G}^{\prime}}(\alpha_{2})=\sigma^{\mathcal{G}}(\alpha_{2})$.
@@
Next, Jedwabny et al. (?) introduced a property which requires that the length of a proof in the logic, i.e., a chain of reasoning, should not affect the strength of a statement.
@@
This property is indeed intuitive in a number of settings where we would prioritise the logical consistency over approximate evaluations of statements. However, in some settings, e.g., when modelling human discourse, it may be the case that the same logical conclusions, when reached over the course of multiple statements, are weaker, or even stronger, than those stated succinctly (e.g., see the study of priority in (?)). For example, given the statements instantiated in rewriting, consider the effect of adding a statement $\alpha_{4}=\langle\Phi_{4},\Psi_{4}\rangle$, where $\Psi\equiv\neg x_{1}$. In some settings it may be desirable that the weakening effect of $\alpha_{4}$ on $\alpha_{3}$ is lesser than that on $\alpha_{1}$ due to the distance. We leave a formalisation of this (non-trivial) property to future work. We would, however, argue that rewriting may not be universally applicable, as illustrated below.
@@
Continuing from Example 2, let us introduce two additional statements in the SG of Figure 1: $\alpha_{5}=\langle a,e\rangle$ (“CO2 emissions are rising, therefore extreme weather events are increasing”), and $\alpha_{6}=\langle e\wedge b,c\rangle$ (“Extreme weather events and global temperatures are increasing, therefore climate change is human-caused”). Let $\tau(\alpha_{5})=1$ and $\tau(\alpha_{6})=0.8$. $\mathcal{A}=\{(\alpha_{4},\alpha_{1}),(\alpha_{4},\alpha_{5})\}$. Under the DC semantics, we get $\sigma_{\wedge_{D}}^{\mathcal{G}}(\alpha_{1})=0.877$ and $\sigma_{\wedge_{D}}^{\mathcal{G}}(\alpha_{6})=\sigma_{D}^{\mathcal{G}^{1}}(e)\times\sigma_{D}^{\mathcal{G}^{1}}(b)=1\times 0.957=0.957$. In contrast, under T-norm-p semantics, we get $\sigma_{T_{p}}^{\mathcal{G}}(\alpha_{1})\!=\!\sigma_{T_{p}}^{\mathcal{G}}(\alpha_{6})\!=\!0.432$. We can see that the DC semantics violate the rewriting property, as $\alpha_{1}$ and $\alpha_{6}$ have different strengths despite making the same claim through different reasoning paths. The T-norm-p semantics, however, is agnostic to the different lengths of the reasoning chain.
Next, Jedwabny et al. (?) state that a statement without at least one supporter for each of its premises, i.e., with no evidence for any of them, is bottom-strength.
@@
This property is clearly useful in cases under complete information, i.e., first case in Definition 9, but in the other two cases, we posit that this property may be too demanding.
@@
In Example 3, weak provability would not require that $\alpha_{4}$’s strength is zero, allowing a form of “trust” in the statement’s weight in the absence of complete information providing the logical reasoning to support its premises. This means the information is not lost in the partially-complete or incomplete cases, unless it is assigned the minimal weight by the GS, which essentially indicates it cannot be trusted.
@@
The desirability of this property is clear: as demonstrated below, it gives an intuitive starting point for the strength of a statement, regardless of its completeness from Definition 9.
@@
First, we consider the effect of adding a statement that has been assigned a zero strength, i.e., it should not affect the strength of any statement it attacks or supports, as in neutrality (?) for abstract GS.